OpenStudy (anonymous):

How would I solve √ 2x+5=7?

OpenStudy (jackellyn):

subtract the five from both sides sqrt(2x)=7-5 square both sides to get ride of the sqrt (sqrt(2x))^2=(7-5)^2 2x=(7-5)^2 Divide by 2 x= ((7-5)^2)/2